OpenStudy (anonymous):

If w varies directly as 3x + 2, and w = 56 when x = 4, find x when w = 92.

OpenStudy (saifoo.khan):

find the value of constant first.

myininaya (myininaya):

w=k(3x+2) 56=k(3*4+2) 56=k(12+2) 56=k(14) -- ---- 14 14 56/14=k k=56/14=4

myininaya (myininaya):

w=k(3x+2) 92=4(3x+2) 92/4=3x+2 23=3x+2 23-2=3x+2-2 21=3x -- -- 3 3 7=x x=7